OpenStudy (anonymous):

A collection of nicels and quarters is worth $3.60. There are 20 coins total. How many of each coin are in the collection?

OpenStudy (anonymous):

n=nickels q=quarters n+q=20 0.05n+0.25q=3.60 5n+25q=360 n+5q=72 n+5(20-n)=72 n+100-5n=72 -4n=-28 n=7 7 nickels 13 quarters